OpenStudy (anonymous):

quadratic system 2x^2 = 4y 2x-2y = -5 please show the work with the answer, i want to know how to do quadratic systems. elimination method if possible, please

sam (.sam.):

\[2x^2 = 4y..............(1)\] \[2x-2y = -5 ...............(2)\] -------------------------------- \[2x^2 = 4y..................(1)\] \[2x= -5+2y ..........(2)\] Square both sides, \[2x^2=4y.............(1)\] \[4x^2=(-5+2y)^2.............(2)\] From (2), \[2x^2=\frac{(2y-5)^2}{2}........(2)\] ------------------------------------ \[2x^2 = 4y..................(1)\] \[2x^2=\frac{(2y-5)^2}{2}........(2)\] Elimination \[0=4y-\frac{(2y-5)^2}{2}\] \[y\text{ = }\frac{1}{2} \left(7\pm2 \sqrt{6}\right)\] Find x using the y valaues As you can see elimination method here is more complicated than substitution method.

OpenStudy (theviper):

@.Sam. we do not need sites as we have a very good helper like u :) Very nice solution :)

OpenStudy (lgbasallote):

it boggles me though @.Sam. ....wouldnt it have been easier to cancel y rather than x? i mean put the equations in the form y =